OpenStudy (anonymous):

What are the solutions of 2x2 + 8x = -26?

OpenStudy (anonymous):

A) -2 ± 3i B)-2 ± i C)2 ± 6i D)-3 ± 2i

OpenStudy (anonymous):

\[2x^2+8x=-26~~~~~~~->~~~~~~~~~x^2+4x=-13\] I divided everything by 2.

OpenStudy (anonymous):

Now, how would you like to do it, by completing the square, quadratic formula? (you can't really factor this)

OpenStudy (anonymous):

how would i complete the square

OpenStudy (anonymous):

You don't have to if you don't never done that before, but if you did and want/have to complete the square, add 4 to both sides.

OpenStudy (anonymous):

if i were to do that would it be x^2 = -4x -13

OpenStudy (anonymous):

Completing the square. \[x^2+4x~=~-13\]\[~~~~~~~~+4~~~~~~~+4\] \[x^2+4x+4=-9~~~~~~~~->~~~~~~~(x+2)^2~=~ (3i)^2\] Can you do it now?

OpenStudy (anonymous):

so lost hah...

OpenStudy (anonymous):

Do you see how I added 4 to both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright, then do you see that \[x^2+4x+4\] factors into \[(x+2)(x+2)\] which is same as \[(x+2)^2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK, so now I am showing a pre-step that would help you to square both sides. I am FINDING the SQUARE ROOT of -9

OpenStudy (anonymous):

\[-9~~= ~~(-1) \times 9 ~~=~~(i^2)~ \times ~3^2~=~(3i)^2\] DO you follow how I found the square root of -9 ?

OpenStudy (anonymous):

um i think so

OpenStudy (anonymous):

Alrighty, but the thing is that the square root can be NEGATIVE OR POSITIVE, okay? So it would be +-3i RIGHT ?

OpenStudy (anonymous):

uhuh.

OpenStudy (anonymous):

Lost?

OpenStudy (anonymous):

no i think I'm starting to get it

OpenStudy (anonymous):

Ok, so see how \[x^2+4x+4=-9~~~~~\] turns into \[(x+2)^2=~~(+-~3i)^2\] which then after squaring both sides, turns into\[x+2=3i~~~~~or~~~~x+2=-3i\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

SO can you preform just the one final step for me? Please tell me your answer and what you did to get it?

OpenStudy (anonymous):

alrigh

OpenStudy (anonymous):

ok i would just subtract 2 from both swides so x=-2 +/- -3 ?

OpenStudy (anonymous):

Yeap! COrrect, x= -2 +- 3i

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

No problem!